OpenStudy (anonymous):

Can someone help me solve this: Find the slope of the functions graph given a point: g(x)=(x)/(x-2) at (3,3). I think i have to use the difference quotient. Thanks

OpenStudy (anonymous):

g(x)=x/x-2 g'=(x-2)*1-x*1 ----------- (x-2)^2 g'= -2/(x-2)^2 slope=-2/1=-2